OpenStudy (anonymous):

find the domain y = x^2-16/x^3+2x^2-9x-18. still not understanding how to do this

OpenStudy (debbieg):

The denominator is \(x^3+2x^2-9x-18\), right?

OpenStudy (debbieg):

What methods have you been taught for finding roots of a 3rd degree polynomial? Graphing? Rational root theorem? Synthetic division?

OpenStudy (anonymous):

The problem is x^2 over x^3+2x^2-9x-18. Looking for the domain.

OpenStudy (anonymous):

sorry the first part is x^2-16 over the rest

OpenStudy (debbieg):

Ok, so like I said - the denominator of the function is \(x^3+2x^2-9x-18\). I just wanted to confirm that since you did not use any ( ) so it can be vague in meaning.

OpenStudy (debbieg):

OK, the numerator isn't really very relevant if we are just looking for the domain. How do you find the domain of a rational function? Or maybe the way to ask it is: what kind of value ISN'T in the domain of a rational function?

OpenStudy (debbieg):

E.g., for the function \(\Large y=\dfrac{ 3 }{ x-2 }\) ..... what is the domain?

OpenStudy (anonymous):

2

OpenStudy (debbieg):

The domain is 2?

OpenStudy (debbieg):

No.... not exactly.....

OpenStudy (anonymous):

everything except 2

OpenStudy (debbieg):

RIGHT. So in general, how do you find the domain of a rational function? It's everything EXCEPT the values that............?

OpenStudy (anonymous):

equal 0

OpenStudy (debbieg):

Well, sort of. It's everything except the X values that make the DENOMINATOR =0. right? Remember - domain is all about the x's. What x's are ALLOWED to be plugged into the equation. When it comes to a rational function, the x's that are NOT allowed are those that make the den'r=0.

OpenStudy (debbieg):

So what you need to do here, if find all the x's that make this den'r=0. That is, you have to solve: \(\Large x^3+2x^2-9x-18=0\) Which brings me back to.... :) What methods have you been taught for finding roots of a 3rd degree polynomial? Graphing? Rational root theorem? Synthetic division?

OpenStudy (debbieg):

When you solve that equation, you'll know what has to be left out of the domain.

OpenStudy (anonymous):

rational root

OpenStudy (debbieg):

OK. so by rational root theorem, any rational root of this function (if one exists) must be a factor of what number?

OpenStudy (anonymous):

2

OpenStudy (debbieg):

That rational root theorem says that IF a number is rational root, then it is of the form \(\pm\)p/q, where p is a factor of the constant term (in this case, that is 18) and q is a factor of the leading coefficient (in this case, 1). so you need to find all factors of p=18, and q=1, (so in this case, really just all factors of 18). THAT list includes any rational roots. Notice that it doesn't guarantee that there are ANY rational roots - but only that, if there ARE, they are on that list.

OpenStudy (anonymous):

ok so i got a domain of everything except for 3,-3 and -2

OpenStudy (debbieg):

Very good - that's it! :)

OpenStudy (anonymous):

Thank you